OpenStudy (anonymous):

How do you find the focus of a parabola with the equation x=4(y-2)^2 +3 ??

OpenStudy (saifoo.khan):

it means vertex right?

OpenStudy (amistre64):

it means focus

OpenStudy (amistre64):

you have to find the vertex yes :) but then measure from it up by "a"

OpenStudy (amistre64):

y^2 = 4ax is the general setup for this where a is the distance we seek

OpenStudy (amistre64):

(x-3) = 4(y-2)^2 x = 4y^2 1/4 x = y^2 4a = 1/4 ; a = 1/8

OpenStudy (amistre64):

since our vertex is at 3,2 ; and we open up to right; is say the focus is at (3+1/8 , 2)

OpenStudy (amistre64):

i think my a is off http://www.wolframalpha.com/input/?i=%28x-3%29+%3D+4%28y-2%29%5E2

OpenStudy (amistre64):

a = 1/16 :)

OpenStudy (amistre64):

thats better; (3+1/16, 2)

OpenStudy (amistre64):

the key tho is to know the geometric eq for the parabola ....

OpenStudy (anonymous):

Thank you! This was really helpful!

OpenStudy (amistre64):

yw